# Lie Group Theory of Multipole Moments and Shape of Stationary Rotating Fluid Bodies

### v3

@@ 1.2.5 The Zharkov-Trubitsyn Framework for Rotating Giant Planets
$V({\bm{y}})$ $=$ $\frac{4\pi G\bar{\rho}(r)r^{2}}{3}\sum_{n=0}A_{2n}(r)P_{2n}(\cos\theta)\;.$ (1.9)
Here, $\bar{\rho}(r)$ is the average density of the fluid within the volume of radius $r$, and the coefficients $A_{2n}(r)$ represent an algebraic sum of integrals involving the fluid density $\rho=\rho(r)$, the radial harmonics $s_{2n}$, and their multiple products.
The condition (1.7) requires that $A_{2n}(r)=0$ $\forall n\geq 2$. This results in a set of integro-algebraic equations for the radial harmonics $s_{2n}(r)$. The coefficient $A_{0}(r)$ fixes the value of the potential $V({\bm{y}})$ in the left-hand side of Eq. (1.7) by setting, $A_{0}(r):=\displaystyle\frac{3}{4\pi G\bar{\rho}(r)r^{2}}V(r)$ on the level surface of the body. This provides a supplementary condition for determining the radial distribution of density $\rho=\rho(r)$ for a given equation of state $p=p(\rho)$ by solving the equation of hydrostatic equilibrium (1.1) on the level surfaces in terms of the radial harmonics $s_{2n}(r)$. Substituting the resulting density $\rho(r)$ into the system of integro-algebraic equations and solving it, the radial harmonics $s_{2n}(r)$ can be determined. The values of the harmonics $s_{2n}(r)$ at the surface of the body are crucial to determine theoretical values of the zonal harmonic coefficients $J_{2n}$ in the multipolar expansion of the body’s gravitational field. Comparison of the theoretical and experimental values of $J_{2n}$ allows to analyze the interior structure of the celestial body [2, 51, 54].
@@ 1.4 Structure and Organization of the Paper
Section 3 discusses the linear theory of infinitesimal perturbations of the rotating fluid caused by the quadrupole component $W_{Q}$ of the centrifugal potential. The theory operates on the base manifold $\mathfrak{M}\in\mathbb{R}^{3}$, which consists of all internal points $\bm{x}$} of the spherically symmetric volume $\mathcal{V}$} of the body reference configuration, including its boundary $\partial{\mathcal{V}}$. Section 3 deals with the Eulerian variations of the fluid density $\delta_{\bm{\xi}}\rho$, pressure $\delta_{\bm{\xi}}p$, and gravitational potential $\delta_{\bm{\xi}}{\cal U}$. These variations are induced by the infinitesimal translation $\xi^{i}=\xi^{i}({\bm{x}})$ – known as a vector diffeomorphisms – of a fluid element from its undisturbed position ${\bm{x}}\in\mathfrak{M}$ in the reference configuration to the displaced point ${\bm{x}}+{\bm{\xi}}\in\mathfrak{M}^{\dagger}$. Here, $\mathfrak{M}^{\dagger}$ denotes the deformed base manifold, corresponding to the three-dimensional volume $\mathcal{V}$$^{\dagger}$ occupied by the fluid after the external perturbation $W_{Q}$ has been applied. Mathematical consideration of the Eulerian variations is most effectively done within the framework of the Lie algebra $\mathfrak{g}$ of diffeomorphisms $\bm{\xi}$}, which operate in the tangent space $T\mathfrak{M}$ of the base manifold $\mathfrak{M}$. The Lie algebra $\mathfrak{g}$ is equipped with the Lie bracket ${\bm{\xi}},{\bm{\eta}}]$ of the vector fields ${\bm{\xi}}\in\mathfrak{g}$ and ${\bm{\eta}}\in\mathfrak{g}$, which naturally coincides with the Lie derivative $\pounds_{\bm{\xi}}{\bm{\eta}}=[{\bm{\xi}},{\bm{\eta}}]$ of one vector field with respect to another. The Lie derivative defines the infinitesimal Eulerian variation of density, $\delta_{\bm{\xi}}\rho\equiv\pounds_{\bm{\xi}}\rho$, and pressure, $\delta_{\bm{\xi}}p\equiv\pounds_{\bm{\xi}}p$, and is also used for the definition of the infinitesimal variation $\delta_{\bm{\xi}}{\mathcal{U}}$ of the gravitational potential $\mathcal{U}$}. We demonstrate that the density variation $\delta_{\bm{\xi}}\rho$ couples linearly with the perturbation of the gravitational field $\delta_{\bm{\xi}}\mathcal{U}$. This linear coupling plays a central role in deriving the Helmholtz equation for the total gravitational field perturbation, defined as ${\cal K}\equiv\delta_{\bm{\xi}}\mathcal{U}+W_{Q}$. This relation is also known as Molodensky’s equation [61]. We further show how the gravitational field perturbation $\cal K$} is connected to the geometry of the level surfaces, and we employ Molodensky’s equation to derive the classical Clairaut equation, which describes the figure of a slowly rotating fluid body.
@@ 2 Base Manifold and Reference Configuration of a Rotating Fluid Body
In the course of this study, we have determined that, from the mathematical standpoint of perturbation theory, it is most optimal to consider the monopole component of the rotational perturbation as an integral part of the undisturbed configuration of the body, which we refer to as the reference configuration. The reference configuration has a fixed radius $r=a$, which depends on the constant angular velocity $\omega$ as a parameter. Non-radial distortions of the body are calculated relative to this reference configuration. This section outlines the equations used in modeling the reference configuration. We adopt a Cartesian coordinate system ${\bm{x}}\epsilon:=(x^{i})=(x^{1},x^{2},x^{3})$, fixed to the rigidly rotating reference configuration of the fluid body. The $x^{3}=z$ axis is aligned with the angular velocity vector $\omega^{i}=(0,0,\omega)$, where $\omega$ denotes the constant angular speed of rotation.
@@ 4.1 Fluid Diffeomorphisms: From Lie Algebra to Lie Group
$w^{i}=x^{i}+{\boldsymbol{\mathscr{X}}}_{\tau}^{i}({\bm{x}})\;,$ (4.4)
where the vector ${\boldsymbol{\mathscr{X}}}_{\tau}^{i}$ represents the finite displacement of the fluid parcel and is given by the exponential map:
${\boldsymbol{\mathscr{X}}}_{\tau}^{i}({\bm{x}})=\left[\exp\left(\tau L_{\bm{\xi}}\right)-1\right]x^{i}=\sum_{n=1}^{\infty}\frac{\tau^{n}}{n!}L_{\bm{\xi}}^{n}x^{i}\;.$ (4.5)
@@
${\boldsymbol{\mathscr{X}}}_{\tau}^{i}({\bm{x}})$ $=$ $\sum_{n=0}^{\infty}\frac{\tau^{n+1}}{(n+1)!}L_{\bm{\xi}}^{n}\xi^{i}\;.$ (4.6)
The vector ${\boldsymbol{\mathscr{X}}}_{\tau}^{i}\in{\mathfrak{G}}_{\tau}$ generalizes the infinitesimal translation $\xi^{i}\in\mathfrak{g}_{\bm{x}}$ introduced in Eq. (3.1) to the domain of a finite size. Here, the coordinates $w^{i}\in{\mathfrak{M}}_{\tau}$ represent the Eulerian position of the fluid element, while $x^{i}\in{\mathfrak{M}}$ are its Lagrangian coordinates on the base manifold $\mathfrak{M}$}.
@@ 4.3.1 Gauge Freedom in Lie Algebra Generators
$\delta_{\bm{\eta}}\rho=-\pounds_{\bm{\eta}}\rho={\bm{\nabla}}\cdot(\rho{\bm{\eta}})=0\;,$ (4.21)
2where ${\bm{\nabla}}=(\partial_{i})$ denotes the gradient operator.
@@
$\delta_{\bm{\eta}}\rho$ $=$ $\pounds_{\bm{\eta}}\rho={\bm{\nabla}}\cdot\left[\rho\left({\bm{\alpha}}\times{\bm{x}}\right)\right]={\bm{\nabla}}\rho\cdot\left({\bm{\alpha}}\times{\bm{x}}\right)-\rho{\bm{\alpha}}\cdot\left({\bm{\nabla}}\times{\bm{x}}\right)\;.$ (4.24)
@@ 4.3.3 The Radial Gauge
The presence of gauge freedom implies that, in the context of hydrostatic perturbations, an identical variation in fluid density within the perturbed configuration can be produced through multiple distinct permutations of fluid parcels from their original positions in the unperturbed reference state. These permutations are governed by the gauge vector field $\bm{\eta}$}, as defined in Eq. (4.22). Although such rearrangements alter the geometry of the level surfaces, this transformation can always be interpreted as resulting from displacements of fluid parcels exclusively along radial directions from their initial positions.
@@ 5.1 Gravitational Potential and Level Surfaces
The rotation-induced change in density at each point on the base manifold $\mathfrak{M}$} is accompanied by a geometric deformation of the manifold itself, ${\mathfrak{M}}\to{\mathfrak{M}}_{\tau}$, and a rearrangement of fluid parcels. This rearrangement alters both the shape of the reference configuration and, in the case of a compressible fluid, its volume $\mathcal{V}$}. Importantly, the vector flow generator $\xi^{i}$, which defines the fluid displacement, does not fully determine the deformation of the base manifold $\mathfrak{M}$}. Likewise, the finite diffeomorphism ${\boldsymbol{\mathscr{X}}}_{\tau}^{i}$ alone does not uniquely define the shape of the rotating body.
@@
In the general case of a compressible fluid, $\zeta({\bm{x}})\neq\xi({\bm{x}})$, and the displacement $X^{i}_{\tau}\neq{\boldsymbol{\mathscr{X}}}^{i}_{\tau}$. These quantities coincide only in the incompressible fluid case. At this stage, explicit expressions for $\zeta$ and $X_{\tau}$ are not required for the discussion of subsequent transformations of the gravitational potential. The generator $\zeta$ is obtained by solving the Clairaut equation (3.49), while the differential equation governing the height function $X_{\tau}$ will be derived in Section 6.3.2.
@@
However, the integration domain ${\cal V}_{\tau}$ in Eq. (5.5) corresponds to the perturbed configuration, mapped from the reference volume $\mathcal{V}$} via the diffeomorphism ${\mathcal{V}}\to{\cal V}_{\tau}$, as defined by Eq. (5.2). Consequently, the Eulerian variation of the gravitational potential cannot be obtained by directly subtracting the unperturbed potential ${\mathcal{U}}({\bm{x}})$ (see Eq. (2.2)) from ${\mathcal{U}}_{\tau}({\bm{x}})$, since the respective integrals are defined over different manifolds: the base manifold $\mathcal{M}$ for ${\mathcal{U}}({\bm{x}})$ and the perturbed manifold $\mathfrak{N}_{\tau}$ for ${\mathcal{U}}_{\tau}({\bm{x}})$.
To address this, and following the approach used in the linearized perturbation theory of Section 3.4, we apply a pullback diffeomorphism to transform the integration in Eq. (5.5) from manifold $\mathfrak{N}_{\tau}$ back to the base manifold $\mathcal{M}$. This reduces the integration domain to the spherical reference volume $\mathcal{V}$}, enabling the decomposition of ${\mathcal{U}}_{\tau}$ into volume and surface integrals.
@@ 5.3 Poisson Equation for Total Gravitational Perturbation
$K_{\tau}({\bm{x}})$ $:=$ $\left[\exp\left(\tau\delta_{\bm{\xi}}\right)-1\right]U({\bm{x}})+\tau W_{Q}({\bm{x}})=\delta_{\bm{\xi}}{\mathcal{U}}({\bm{x}})+\frac{1}{2}\delta^{2}_{\bm{\xi}}{\mathcal{U}}({\bm{x}})+\ldots+\tau W_{Q}({\bm{x}})\;.$ (5.27)
@@
$\nu({\bm{x}})$ $=$ $\sum_{n=0}^{\infty}\frac{X^{n+1}_{\tau}({\bm{a}})}{(n+1)!}\frac{\partial^{n}}{\partial a^{n}}\left[\frac{a^{2}}{r^{2}}\rho_{\tau}({\bm{a}})\delta(r-a)\right]\;.$ (5.29)
@@ 5.4 Nonlinear Interaction between Fluid Density and Gravitational Field Perturbations
$\delta_{\bm{\xi}}U+\tau W_{Q}$ $=$ ${\cal A}\delta_{\bm{\xi}}\rho\;,$ (5.31)
where we use a $\tau$-parameterized definition of $\cal K$} from Eq. (3.29) and introduce the shorthand notation:
@@ 6.1.1 Functional Equation for Level Surfaces in Rotating Fluid Bodies
We introduce the notation fro the effective gravitational potential $U(r(\rho)):=\mathsf{U}(\rho)$ and express Eq. (5.46) in terms of $\mathsf{U}$:

### v4

@@ 1.2.5 The Zharkov-Trubitsyn Framework for Rotating Giant Planets
$V({\bm{y}})$ $=$ $\frac{4\pi Gr^{2}}{3}\sum_{n=0}A_{2n}(r)P_{2n}(\cos\theta)\;.$ (1.9)
Here, the coefficients $A_{2n}(r)$ represent an algebraic sum of integrals involving the fluid density $\rho=\rho(r)$, the radial harmonics $s_{2n}$, and their multiple products.
The condition (1.7) requires that $A_{2n}(r)=0$ for all $n\geq 1$. This results in a set of integro-algebraic equations for the radial harmonics $s_{2n}(r)$. The coefficient $A_{0}(r)$ fixes the value of the potential $V({\bm{y}})$ in the left-hand side of Eq. (1.7) by setting, $A_{0}(r):=\displaystyle\frac{3}{4\pi Gr^{2}}V(r)$ on the level surface of the body. This provides a supplementary condition for determining the radial distribution of density $\rho=\rho(r)$ for a given equation of state $p=p(\rho)$ by solving the equation of hydrostatic equilibrium (1.1) on the level surfaces in terms of the radial harmonics $s_{2n}(r)$. Substituting the resulting density $\rho(r)$ into the system of integro-algebraic equations and solving it, the radial harmonics $s_{2n}(r)$ can be determined. The values of the harmonics $s_{2n}(r)$ at the surface of the body are crucial to determine theoretical values of the zonal harmonic coefficients $J_{2n}$ in the multipolar expansion of the body’s gravitational field. Comparison of the theoretical and experimental values of $J_{2n}$ allows to analyze the interior structure of the celestial body [2, 51, 54].
@@ 1.4 Structure and Organization of the Paper
Section 3 discusses the linear theory of infinitesimal perturbations of the rotating fluid caused by the quadrupole component $W_{Q}$ of the centrifugal potential. The theory operates on the base manifold $\mathfrak{M}\in\mathbb{R}^{3}$, which consists of all internal points $\bm{x}$} of the spherically symmetric volume $\mathcal{V}$} of the body reference configuration, including its boundary $\partial{\mathcal{V}}$. Section 3 deals with the Eulerian variations of the fluid density $\delta_{\bm{\xi}}\rho$, pressure $\delta_{\bm{\xi}}p$, and gravitational potential $\delta_{\bm{\xi}}{\cal U}$. These variations are induced by the infinitesimal translation $\xi^{i}=\xi^{i}({\bm{x}})$ – known as a vector diffeomorphism – of a fluid element from its undisturbed position ${\bm{x}}\in\mathfrak{M}$ in the reference configuration to the displaced point ${\bm{x}}+{\bm{\xi}}\in\mathfrak{M}^{\dagger}$. Here, $\mathfrak{M}^{\dagger}$ denotes the deformed base manifold, corresponding to the three-dimensional volume $\mathcal{V}$$^{\dagger}$ occupied by the fluid after the external perturbation $W_{Q}$ has been applied. Mathematical consideration of the Eulerian variations is most effectively done within the framework of the Lie algebra $\mathfrak{g}$ of diffeomorphisms $\bm{\xi}$}, which operate in the tangent space $T\mathfrak{M}$ of the base manifold $\mathfrak{M}$. The Lie algebra $\mathfrak{g}$ is equipped with the Lie bracket ${\bm{\xi}},{\bm{\eta}}]$ of the vector fields ${\bm{\xi}}\in\mathfrak{g}$ and ${\bm{\eta}}\in\mathfrak{g}$, which naturally coincides with the Lie derivative $\pounds_{\bm{\xi}}{\bm{\eta}}=[{\bm{\xi}},{\bm{\eta}}]$ of one vector field with respect to another. The Lie derivative defines the infinitesimal Eulerian variation of density, $\delta_{\bm{\xi}}\rho\equiv\pounds_{\bm{\xi}}\rho$, and pressure, $\delta_{\bm{\xi}}p\equiv\pounds_{\bm{\xi}}p$, and is also used for the definition of the infinitesimal variation $\delta_{\bm{\xi}}{\mathcal{U}}$ of the gravitational potential $\mathcal{U}$}. We demonstrate that the density variation $\delta_{\bm{\xi}}\rho$ couples linearly with the perturbation of the gravitational field $\delta_{\bm{\xi}}\mathcal{U}$. This linear coupling plays a central role in deriving the Helmholtz equation for the total gravitational field perturbation, defined as ${\cal K}\equiv\delta_{\bm{\xi}}\mathcal{U}+W_{Q}$. This relation is also known as Molodensky’s equation [61]. We further show how the gravitational field perturbation $\cal K$} is connected to the geometry of the level surfaces, and we employ Molodensky’s equation to derive the classical Clairaut equation, which describes the figure of a slowly rotating fluid body.
@@ 2 Base Manifold and Reference Configuration of a Rotating Fluid Body
In the course of this study, we have determined that, from the mathematical standpoint of perturbation theory, it is most optimal to consider the monopole component of the rotational perturbation as an integral part of the undisturbed configuration of the body, which we refer to as the reference configuration. The reference configuration has a fixed radius $r=a$, which depends on the constant angular velocity $\omega$ as a parameter. Non-radial distortions of the body are calculated relative to this reference configuration. This section outlines the equations used in modeling the reference configuration. We adopt a Cartesian coordinate system ${\bm{x}}:=(x^{i})=(x^{1},x^{2},x^{3})$, fixed to the rigidly rotating reference configuration of the fluid body. The $x^{3}=z$ axis is aligned with the angular velocity vector $\omega^{i}=(0,0,\omega)$, where $\omega$ denotes the constant angular speed of rotation.
@@ 4.1 Fluid Diffeomorphisms: From Lie Algebra to Lie Group
$w^{i}=x^{i}+{\mathscr{X}}_{\tau}^{i}({\bm{x}})\;,$ (4.4)
where the vector ${\boldsymbol{\mathscr{X}}}_{\tau}=({\mathscr{X}}_{\tau}^{i})$ represents the finite displacement of the fluid parcel and is given by the exponential map:
${\mathscr{X}}_{\tau}^{i}({\bm{x}})=\left[\exp\left(\tau L_{\bm{\xi}}\right)-1\right]x^{i}=\sum_{n=1}^{\infty}\frac{\tau^{n}}{n!}L_{\bm{\xi}}^{n}x^{i}\;.$ (4.5)
@@
${\mathscr{X}}_{\tau}^{i}({\bm{x}})$ $=$ $\sum_{n=0}^{\infty}\frac{\tau^{n+1}}{(n+1)!}L_{\bm{\xi}}^{n}\xi^{i}\;.$ (4.6)
The vector ${\mathscr{X}}_{\tau}^{i}\in{\mathfrak{G}}_{\tau}$ generalizes the infinitesimal translation $\xi^{i}\in\mathfrak{g}_{\bm{x}}$ introduced in Eq. (3.1) to the domain of a finite size. Here, the coordinates $w^{i}\in{\mathfrak{M}}_{\tau}$ represent the Eulerian position of the fluid element, while $x^{i}\in{\mathfrak{M}}$ are its Lagrangian coordinates on the base manifold $\mathfrak{M}$}.
@@ 4.3.1 Gauge Freedom in Lie Algebra Generators
$\delta_{\bm{\eta}}\rho=\pounds_{\bm{\eta}}\rho=-{\bm{\nabla}}\cdot(\rho{\bm{\eta}})=0\;,$ (4.21)
where ${\bm{\nabla}}=(\partial_{i})$ denotes the gradient operator.
@@
$\delta_{\bm{\eta}}\rho$ $=$ $\pounds_{\bm{\eta}}\rho=-{\bm{\nabla}}\cdot\left[\rho\left({\bm{\alpha}}\times{\bm{x}}\right)\right]=-{\bm{\nabla}}\rho\cdot\left({\bm{\alpha}}\times{\bm{x}}\right)+\rho{\bm{\alpha}}\cdot\left({\bm{\nabla}}\times{\bm{x}}\right)\;.$ (4.24)
@@ 4.3.3 The Radial Gauge
The presence of gauge freedom implies that, in the context of hydrostatic perturbations, an identical variation in fluid density within the perturbed configuration can be produced through multiple distinct permutations of fluid parcels from their original positions in the unperturbed reference state. These permutations are governed by the gauge vector field $\bm{\eta}$}, as defined in Eq. (4.22). Although such rearrangements can be implemented in various ways, they do not alter the geometry of the level surfaces and can always be interpreted as resulting from displacements of fluid parcels exclusively along radial directions from their initial positions.
@@ 5.1 Gravitational Potential and Level Surfaces
The rotation-induced change in density at each point on the base manifold $\mathfrak{M}$} is accompanied by a geometric deformation of the manifold itself, ${\mathfrak{M}}\to{\mathfrak{M}}_{\tau}$, and a rearrangement of fluid parcels. This rearrangement alters both the shape of the reference configuration and, in the case of a compressible fluid, its volume $\mathcal{V}$}. Importantly, the vector flow generator $\xi^{i}$, which defines the fluid displacement, does not fully determine the deformation of the base manifold $\mathfrak{M}$}. Likewise, the finite diffeomorphism ${\mathscr{X}}_{\tau}^{i}$ alone does not uniquely define the shape of the rotating body.
@@
In the general case of a compressible fluid, $\zeta({\bm{x}})\neq\xi({\bm{x}})$, and the displacement $X^{i}_{\tau}\neq{\mathscr{X}}^{i}_{\tau}$. These quantities coincide only in the incompressible fluid case. At this stage, explicit expressions for $\zeta$ and $X_{\tau}$ are not required for the discussion of subsequent transformations of the gravitational potential. The generator $\zeta$ is obtained by solving the Clairaut equation (3.49), while the differential equation governing the height function $X_{\tau}$ will be derived in Section 6.3.2.
@@
However, the integration domain ${\cal V}_{\tau}$ in Eq. (5.5) corresponds to the perturbed configuration, mapped from the reference volume $\mathcal{V}$} via the diffeomorphism ${\mathcal{V}}\to{\cal V}_{\tau}$, as defined by Eq. (5.2). Consequently, the Eulerian variation of the gravitational potential cannot be obtained by directly subtracting the unperturbed potential ${\mathcal{U}}({\bm{x}})$ (see Eq. (2.2)) from ${\mathcal{U}}_{\tau}({\bm{x}})$, since the respective integrals are defined over different manifolds: the base manifold $\mathfrak{M}$ for ${\mathcal{U}}({\bm{x}})$ and the perturbed manifold $\mathfrak{N}_{\tau}$ for ${\mathcal{U}}_{\tau}({\bm{x}})$.
To address this, and following the approach used in the linearized perturbation theory of Section 3.4, we apply a pullback diffeomorphism to transform the integration in Eq. (5.5) from manifold $\mathfrak{N}_{\tau}$ back to the base manifold $\mathfrak{M}$. This reduces the integration domain to the spherical reference volume $\mathcal{V}$}, enabling the decomposition of ${\mathcal{U}}_{\tau}$ into volume and surface integrals.
@@ 5.3 Poisson Equation for Total Gravitational Perturbation
$K_{\tau}({\bm{x}})$ $:=$ $\left[\exp\left(\tau\delta_{\bm{\xi}}\right)-1\right]U({\bm{x}})+\tau W_{Q}({\bm{x}})=\tau\delta_{\bm{\xi}}{\mathcal{U}}({\bm{x}})+\frac{\tau^{2}}{2}\delta^{2}_{\bm{\xi}}{\mathcal{U}}({\bm{x}})+\ldots+\tau W_{Q}({\bm{x}})\;.$ (5.27)
@@
$\nu_{\tau}({\bm{x}})$ $=$ $\sum_{n=0}^{\infty}\frac{X^{n+1}_{\tau}({\bm{a}})}{(n+1)!}\frac{\partial^{n}}{\partial a^{n}}\left[\frac{a^{2}}{r^{2}}\rho_{\tau}({\bm{a}})\delta(r-a)\right]\;.$ (5.29)
@@ 5.4 Nonlinear Interaction between Fluid Density and Gravitational Field Perturbations
$\delta_{\bm{\xi}}U+W_{Q}$ $=$ ${\cal A}\delta_{\bm{\xi}}\rho\;,$ (5.31)
where we use a definition of $\cal K$} from Eq. (3.29) and introduce the shorthand notation:
@@ 6.1.1 Functional Equation for Level Surfaces in Rotating Fluid Bodies
We introduce the notation for the effective gravitational potential $U(r(\rho)):=\mathsf{U}(\rho)$ and express Eq. (5.46) in terms of $\mathsf{U}$: